# MicroScopiQ: Accelerating Foundational Models through Outlier-Aware Microscaling Quantization

### v3

@@ 6.2. LLM Quantization Results
Weight-Only Quantization. In Table 4, we compare the WikiText2 PPL of different LLMs for W4A16 and W2A16 settings. MicroScopiQ consistently outperforms all the baselines across different models and quantization settings. At W4A16, MicroScopiQ achieves near-lossless quantization performance. Particularly at W2A16, the benefits of the MicroScopiQ method is evident, achieving up to a 2.04 decrease in PPL compared to the baselines. This is due to MicroScopiQ’s ability to quantize outliers at higher-precision and in FP-format to reduce quantization error. At W2A16, techniques like AWQ and OliVe fail ($\geq 1e4$ PPL, not depicted in table). MicroScopiQ outperforms the SoTA training-based framework OmniQuant across all models which requires 56×5-6$\times$5 - 6 × higher runtime.
Weight-Activation Quantization. Similarly, we compare the quantization performance of MicroScopiQ for W4A4 and W2A8 configurations in Table 4. Due to the dynamic nature of activations, very rarely do techniques quantize outliers in activation directly (Zhao et al., 2024). Instead, most techniques (Xiao et al., 2023; Shao et al., 2023) migrate the activation outliers to weights to enable simpler activation quantization. We borrow the migration strength ($\alpha$) hyper-parameter introduced in SmoothQuant (Xiao et al., 2023) to migrate the difficulty of quantizing outliers in activations to weights. We find that prior techniques (Xiao et al., 2023; Shao et al., 2023) can only migrate up to 50% of activation quantization difficulty to weights, i.e., $\alpha=0.5$ before beginning to introduce higher weight quantization error. MicroScopiQ’s robustness to higher presence of outliers in weights by effectively quantizing outliers at higher precision and identifying least important weights to prune, allows $\alpha$ as high as 0.7 i.e., migrating most of the activation outliers to weights. This enables simpler activation quantization with MX-INT-(4/8)128 and absolving the need to handle outliers in activations. With $\alpha=0.7$ for MicroScopiQ, $\alpha=0.5$ for SmoothQuant and learned migration strength for OmniQuant in Table 4, MicroScopiQ outperforms all baselines with up to 7.4$\times$ lower PPL across quantization settings. Furthermore, MicroScopiQ outperforms a recent work Atom (Zhao et al., 2024) that performs activation quantization with up to a 0.33 lower PPL.
Figure 8. Weight-only quantization for VLMs, OpenFlamingo-9B (Awadalla et al., 2023) on COCO and VILA-7B (Lin et al., 2024b) on VizWiz and TextVQA.
@@ 6.4. Accelerator Results
Compute Area. In Table 5, we compare the accelerator compute area breakdown of MicroScopiQ (w/ 1 ReCoN unit) with baselines GOBO and OliVe for a $64\times 64$ array design. For a fair evaluation, all accelerators have identical configurations with same number of PEs. MicroScopiQ has a very low compute overhead $8.63\%$, compared to OliVe’s $9.90\%$. This is because, units like ReCoN which perform three simple operations for outlier processing has a very low area utilization compared to OliVe’s encoders and decoders. ReCoN is also time-multiplexed across all rows resulting in minimal area overhead (see section 6.5). Furthermore. the simple INT operations of MicroScopiQ allows the packing of more compute power with minimal overheads. A similar multi-precision design in OliVe would require 3.5-4$\times$ higher multi-precision support area due to complex exponent-integer PEs. GOBO has an order of magnitude higher compute area compared to OliVe and MicroScopiQ due to large per-PE area.
@@
Maximum Performance Per Unit Area ($TOPS/mm^{2}$). We also estimate the peak compute throughput per unit area of each accelerator, using LLaMA3-8B as the workload and report it as the compute density in Table 5. To achieve peak throughput for MicroScopiQ, FMs must be quantized to low precision (e.g.. $bb=2$). MicroScopiQ achieves nearly 2$\times$ and 14$\times$ higher compute density than OliVe and GOBO respectively.
@@
Iso-Accuracy Energy Comparison. In Figure 9(c), we show the normalized energy consumption of different accelerators, composed of static and dynamic energy. MicroScopiQ v2 has the lowest energy consumption. Compared to designs like GOBO, OLAccel and AdaptivFloat, where computations happen at higher-precisions in 8-bit and 32-bit PEs, MicroScopiQ v2 has a significant advantage in terms of the core and DRAM dynamic energy consumption with simple INT PEs. On average MicroScopiQ v2, has 1.5$\times$ lower energy consumption compared to all baselines across different FMs.
@@
$+$ 2-bit KV-cache quantization (Liu et al., 2024b) 9.58 ($\uparrow 0.50$)
@@ 6.5. Algorithm ablations
KV-cache quantization. In Table 6 we also report the impact of KV-cache quantization on performance. Following the 2-bit KV-cache quantization technique proposed in (Liu et al., 2024b) we quantize K per channel, and the V per token with a MaB of size 128 and residual token length (R) (Liu et al., 2024b) of 128 for both key and value.
@@
Figure 10. (a) Effect of outlier group size on the Wikitext-v2 perplexity (PPL) and EBW of a MicroScopiQ quantized LLaMA3-8B. $\infty$ is per-tensor granularity. (b) Percentage access conflicts to ReCoN for $64\times 64$ PE array.
Outlier group size. In Figure 10(a), we analyze the effect of different $B_{\mu}$ values on outlier diversity (measure by the standard deviation: red line), PPL (green bar), and EBW (blue line) for LLaMa3-8B. For $B_{\mu}=2,4$, accuracy degradation occurs due to outlier pruning, as many mBs have $\geq B_{\mu}/2$ outliers. As $B_{\mu}$ increases, outliers further apart show greater diversity, evidenced by the higher standard deviation. For $B_{\mu}\geq 32$, this results in increased quantization error and higher PPL, as MX-FP scales are shared across more diverse outlier values. Larger group sizes also significantly increase EBW due to metadata overhead. A balance is achieved at $B_{\mu}=8$.
Figure 11. (a) Different variants of MicroScopiQ accelerator with varying number of ReCoN units. (b) Annotated MicroScopiQ floorplan ($64\times 64$ array).
@@ 6.6. Analysis of time-multiplexed ReCoN
Accelerator design variants. InFigure 11(a), we show the layout of different variants of MicroScopiQ with varying number of ReCoN units, each of which can be employed based on application requirements. For area critical applications, design A, with a single ReCoN can be employed to have minimal area utilization with a slight increase in latency due to access conflicts to ReCoN. With the addition of more ReCoN units, fewer rows of PEs share a single ReCoN, dramatically reducing access conflicts to the shared ReCoN unit (design B). When the number of ReCoN units equals the number of rows, each PE row gets a dedicated ReCoN unit as shown in design C, which can be employed for latency critical applications. As we shall demonstrate subsequently, we identified that in up to a $128\times 128$ PE array size, a total of 8 ReCoN units are sufficient to achieve peak performance with zero access conflicts to ReCoN.
Optimal # of ReCoN units. For a $64\times 64$ MicroScopiQ accelerator, we study the number of access conflicts for different number of ReCoN units in Figure 10(b). Evidently, the number of access conflicts to ReCoN which is measured as the percent of total number of accesses to ReCoN that result in conflicts is under 3%. With the progressive addition of more ReCoN units, the access conflicts tend to 0%. In Figure 9, despite the access conflicts, 1 ReCoN unit is sufficient to reach good inference performance in an iso-accuracy scenario. For latency critical applications, in Figure 13(a), we quantify the performance improvement and impact on compute area by incorporating multiple ReCoN units for a LLaMA3-8B FM. Up to 8 ReCoN units enables peak performance with 21% improved latency and only 1.58×1.58$\times$1.58 × higher compute area. For large array sizes, up to 8 ReCoN units are sufficient for peak performance.
@@ 6.7. Scalability and Timing Analysis
Scalability. To study how MicroScopiQ scales to different PE array sizes, we compare the total area (including on-chip weight and iAct/oAct buffers) against OliVe at three different scales: $8\times 8,16\times 16,128\times 128$ in Figure 12. Following (Sumbul et al., 2022), for $8\times 8$ array we employ 16 kB iAct and oAct buffers and 32 kB weight buffer. We progressively scale the buffers for larger arrays to ensure full array utilization. Since varying the number of ReCoN units trades off area for performance, we also compare three versions of MicroScopiQ with 1, 2, 8 ReCoN units. For the single ReCoN variant of MicroScopiQ, as the design scales to higher array sizes, the area overhead of ReCoN decreases ($128\times 128$ has overhead of 3%), with the on-chip area being dominated by the large PE array and buffers. We observe similar trends for 2 or 8 ReCoN units. At large PE array sizes such as $128\times 128$, adding 8 ReCoN units results in a minimal overhead of just 11%. Particularly at large array sizes the column-wise bus from PEs to ReCoN can be routed over logic as the PE array scales. With same buffer configurations, OliVe has a higher area compared to a single ReCoN MicroScopiQ variant, while comparable area to MicroScopiQ with 8 ReCoN units. At all variants, MicroScopiQ has 23×2-3$\times$2 - 3 × higher performance than OliVe (see Table 5).
Timing Analysis. MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. The critical path for MicroScopiQ is from the local weights registers to multipliers within the PE. The column-wise bus (the longest wire in MicroScopiQ) from PEs to ReCoN is not on the critical path and therefore does not limit scalability  [^6] For extremely large arrays, if global wire becomes a bottleneck, it can be resolved by pipelining/adding more ReCoN units (for timing) or laying wires over logic (for area).. Implementation overhead on NoC-based accelerators. The global column-wise bus is a design commonly seen in AI ASICs like (Chen et al., 2019; Firoozshahian et al., 2023; Shao et al., 2019; Tong et al., 2024). These global buses are used in conjunction with the NoC in these designs for data distribution, layout reordering (Tong et al., 2024). We modify two large-scale accelerators MTIA and Eyeriss-v2 that employ NoCs in their design to support MicroScopiQ, by incorporating ReCoN functionality in the accelerator NoCs and modifying PEs to support all MicroScopiQ specific operations. As shown in Figure 13(b), compared to the baseline design of MTIA and Eyeriss-v2, integrating ReCoN and MicroScopiQ PEs results in only a $3\%$ and $2.3\%$ increase in compute area, respectively. Since these designs already employ NoCs, integration overhead of MicroScopiQ is minimal. This demonstrates that the benefits achieved by MicroScopiQ are attained at minimal to no additional cost.
Figure 13. (a) Effect of time-multiplexed ReCoN units on the compute area and inference latency, (b) Implementation overhead of MicroScopiQ in NoC-based accelerators (Firoozshahian et al., 2023; Chen et al., 2019)

### v4

@@ 7.2. LLM Quantization Results
Weight-Only Quantization. In Table 2, we compare the WikiText2 PPL of different LLMs for W4A16 and W2A16 settings. MicroScopiQ consistently outperforms all the baselines across different models and quantization settings. At W4A16, MicroScopiQ achieves near-lossless quantization performance. Particularly at W2A16, the benefits of the MicroScopiQ method is evident, achieving up to a 2.04 decrease in PPL score compared to the baselines. This is due to MicroScopiQ’s ability to quantize outliers at higher-precision and in FP-format to reduce quantization error. At W2A16, techniques like AWQ and OliVe have unacceptable accuracy degradations ($\geq 1e4$ PPL, not depicted in table). MicroScopiQ outperforms the contemporary SDQ (Jeong et al., 2023) at W2A16 with up to a 2.06 lower PPL score, owing to its flexible pruning+quantization strategy. In contrast, SDQ relies on a rigid N:M pattern, limiting its adaptability across model families. MicroScopiQ also outperforms the SoTA quantization framework OmniQuant across all models which requires 56×5-6$\times$5 - 6 × higher runtime than MicroScopiQ.
Figure 10. Weight-only quantization for VLMs, OpenFlamingo-9B (Awadalla et al., 2023) on COCO, VQAv2 and VILA-7B (Lin et al., 2024c) on VizWiz and TextVQA.
Figure 11. Qualitative results of OpenFlamingo-9B (Awadalla et al., 2023) on 8-shot COCO captioning with weight-only quantization using OliVe (W4) and MicroScopiQ (W2). Text in red and green indicate incorrect and accurate captioning respectively.
Weight-Activation Quantization. Similarly, we compare the quantization performance of MicroScopiQ for W4A4 and W2A8 configurations in Table 2. Due to the dynamic nature of activations, very rarely do techniques quantize outliers in activations directly (Zhao et al., 2024). Instead, most techniques (Xiao et al., 2023; Shao et al., 2023) migrate the activation outliers to weights to enable simpler activation quantization. We borrow the migration strength ($\alpha$) hyper-parameter introduced in SmoothQuant (Xiao et al., 2023) to migrate the difficulty of quantizing outliers in activations to weights. We find that prior techniques (Xiao et al., 2023; Shao et al., 2023) can only migrate up to 50% of activation quantization difficulty to weights, i.e., $\alpha=0.5$ before beginning to introduce higher weight quantization error. MicroScopiQ’s robustness to higher presence of outliers in weights by effectively quantizing outliers at higher precision and identifying least important weights to prune, allows $\alpha$ as high as 0.7 i.e., migrating most of the activation outliers to weights. This enables simpler activation quantization with MX-INT-(4/8)128 and absolving the need to handle outliers in activations. With $\alpha=0.7$ for MicroScopiQ, $\alpha=0.5$ for SmoothQuant and learned migration strength for OmniQuant in Table 2, MicroScopiQ outperforms all baselines with up to 7.4$\times$ lower PPL score across quantization settings. Furthermore, MicroScopiQ outperforms a recent work Atom (Zhao et al., 2024) that performs activation quantization with up to a 0.33 lower PPL score.
@@
Table 4. Quantization results for non-transformer models. We report Top-1 accuracy (%) on ImageNet (higher is better).
  Method W/A ResNet50 VGG16 VMamba-S Vim-S
  Baseline 16/16 76.15% 71.59% 83.60% 80.50%
  HAWQ (Dong et al., 2019) 2/4 73.17% 68.81% - -
QMamba (Li et al., 2025) 4/4 - - 40.36% 68.50%
MicroScopiQ 4/4 75.08% 70.84% 70.07% 71.52%
MicroScopiQ 2/8 75.12% 70.87% 66.52% 71.98%
MicroScopiQ 2/4 73.61% 69.12% - -
@@ 7.3. VLM Quantization Results
We also quantitatively evaluate the performance of MicroScopiQ W2A16 quantized OpenFlamingo-9B VLM on the quality of generated captions for images drawn from the COCO captioning task in Figure 11. The MicroScopiQ quantized model generates accurate captions and preserves overall word semantics compared to OliVe. For instance, in the second figure, OliVe mislabels boat as a van, whereas MicroScopiQ is able to correctly identify it as a boat.
7.4. CNN and SSM Quantization Results
In Table 4, we demonstrate MicroScopiQ quantization on CNNs and SSMs. MicroScopiQ achieves near lossless performance at W4A4, W4A8 configurations and $\leq 3\%$ accuracy drop at W2A4 setting. Similarly, MicroScopiQ achieves up to $30\%$ higher accuracy compared to the SoTA, QMamba (Li et al., 2025), across SSM models.
@@ 7.5. Accelerator Results
Compute Area. In Table 5, we compare the accelerator compute area breakdown of MicroScopiQ (w/ 1 ReCoN unit) with baselines GOBO and OliVe for a $64\times 64$ array design. For a fair evaluation, all accelerators have identical configurations with same number of PEs. MicroScopiQ has a very low compute overhead $8.63\%$, compared to OliVe’s $9.90\%$. This is because, units like ReCoN which perform three simple operations for outlier processing has a very low area utilization compared to OliVe’s encoders and decoders. ReCoN is also time-multiplexed across all rows resulting in minimal area overhead (see section 7.7). Furthermore. the simple INT operations of MicroScopiQ allows the packing of more compute power with minimal overheads. A similar multi-precision design in OliVe would require 3.5-4$\times$ higher multi-precision support area due to complex exponent-integer PEs. GOBO has higher compute area compared to OliVe and MicroScopiQ due to large per-PE area.
@@
Maximum Performance Per Unit Area ($TOPS/mm^{2}$). We also estimate the peak compute throughput per unit area of each accelerator (compute density), using LLaMA3-8B as the workload in Table 5. To achieve peak throughput for MicroScopiQ, FMs must be quantized at $bb=2$. MicroScopiQ achieves nearly 2$\times$ and 14$\times$ higher compute density than OliVe and GOBO respectively.
@@
Iso-Accuracy Energy Comparison. In Figure 12(c), we show the normalized energy consumption of different accelerators, composed of static and dynamic energy. MicroScopiQ v2 has the lowest energy consumption. Compared to designs like GOBO, OLAccel and AdaptivFloat, where computations happen at higher-precisions in 8-bit and 32-bit PEs, MicroScopiQ v2 has a significant advantage in terms of the core and DRAM dynamic energy consumption with simple INT PEs. On average MicroScopiQ v2, has 1.5$\times$ lower energy consumption compared to baselines across different FMs.
Power Breakdown. The power breakdown of MicroScopiQ varies depending on the outlier distribution. For a LLaMA2-7B, the PE array consumes 56.23% of power, followed by on-chip memory (36.80%) and ReCoN (5.94%). Conversely, VILA-7B, characterized by a higher outlier percentage, exhibits increased power consumption in the ReCoN unit (7.65%) and slightly reduced on-chip memory (35.32%), PE array (55.98%) power consumption.
Figure 13. Comparison of MicroScopiQ accelerators with A100 GPU: (a) Normalized latency, (b) Normalized energy.
7.6. GPU Results
Throughput on real GPU. In Table 6, the unoptimized W4A4 MicroScopiQ (MS no-optim) underperforms the FP16 baseline due to the overhead of outlier merging in shared memory and GEMM execution in FP16. In contrast, our optimized W4A4 kernel (MS optim.) (section 6.1), achieves similar performance to SoTA technique Atom (Zhao et al., 2024) due to efficient register caching and dynamic GEMM acceleration, while simultaneously possessing better accuracy.
MicroScopiQ GPU simulation. As shown in our simulated A100 GPU results (Table 6), our tensor core modification enables low-precision INT+FP GEMM, eliminating costly dequantization and FP16 GEMMs, achieving the highest throughput.
Table 6. Normalized token generation throughput comparison across different quantization methods (Wikitext perplexity in brackets) on A100 GPU and GPU simulation (Bakhoda et al., 2009).
  Device Method LLaMA-2 13B LLaMA-3 8B
  A100 TRT-LLM FP16 1.00 (4.83) 1.00 (6.13)
W4A4 Atom 2.25 (6.12) 1.05 (8.12)
W4A4 MS no-optim. 0.98 (5.57) 0.92 (8.12)
W4A4 MS optim. 2.06 (5.57) 1.01 (8.12)
  GPGPU Sim W4A4 MS w/ New MTC 4.31 (5.57) 1.78 (8.12)
      MS: MicroScopiQ; MTC: Modified Tensor Core
GPU v/s MicroScopiQ Accelerator. We compare a standard A100 GPU with the MicroScopiQ accelerator under iso-bandwidth (off-chip: 2 TB/s, on-chip: 1.5kb/cycle/warp) (Raihan et al., 2019) and iso-compute (55,296 multipliers (NVIDIA, 2025)) scenario. As shown in Figure 13(a), MicroScopiQ v1 (W4A4) and v2 (WxA4) (see above) yield 1.2$\times$ and 1.7$\times$ speedup, respectively, over A100 (W4A4), due to multi-precision PEs, avoiding FP16 computation. Additionally, ReCoN’s reordering is pipelined with partial sum reduction, significantly outperforming GPU shfl_sync. Additionally, MicroScopiQ delivers superior energy efficiency (Figure 13(b)), whereas A100 incurs higher on-chip energy consumption, due to register-level reordering and FP16 overhead. Notably, traditional GPUs lack WmAn ($m\neq n$) acceleration, making MicroScopiQ accelerator an efficient alternative.
@@ 7.6. GPU Results
$+$ 2-bit KV-cache quantization (Liu et al., 2024c) 9.58 ($\uparrow 0.50$)
Table 8. LLM quantization performance of MicroScopiQ integrated with OmniQuant.
  Method W/A Llama-2 13B Llama-3 70B Phi-3 3.8B
Baseline 16/16 4.83 2.85 6.33
  OmniQuant (Shao et al., 2023) 4/16 5.02 3.46 6.67
Omni-MicroScopiQ 4/16 4.87 2.97 6.52
  OmniQuant (Shao et al., 2023) 2/16 7.56 6.17 7.09
Omni-MicroScopiQ 2/16 6.58 5.09 6.89
  OmniQuant (Shao et al., 2023) 2/8 8.92 6.83 7.95
Omni-MicroScopiQ 2/8 7.12 5.74 7.21
@@ 7.7. Algorithm ablations
KV-cache quantization. In Table 7 we also report the impact of KV-cache quantization on performance. Following the 2-bit KV-cache quantization technique proposed in (Liu et al., 2024c) we quantize K per channel, and the V per token with a MaB of size 128 and residual token length (R) (Liu et al., 2024c) of 128 for both key and value.
@@
Figure 14. Effect of outlier group size on the Wikitext-v2 perplexity (PPL) and EBW of a MicroScopiQ quantized LLaMA3-8B. $\infty$ is per-tensor granularity.
Outlier group size. In Figure 14, we analyze the effect of different $B_{\mu}$ values on outlier diversity (measure by the standard deviation: red line), PPL (green bar), and EBW (blue line) for LLaMa3-8B. For $B_{\mu}=2,4$, accuracy degradation occurs due to outlier pruning, as many mBs have $\geq B_{\mu}/2$ outliers. As $B_{\mu}$ increases, outliers further apart show greater diversity, evidenced by the higher standard deviation. For $B_{\mu}\geq 32$, this results in increased quantization error and higher PPL, as MX-FP scales are shared across more diverse outlier values. Larger group sizes also significantly increase EBW due to metadata overhead. A balance is achieved at $B_{\mu}=8$.
Figure 15. Different variants of MicroScopiQ accelerator with varying number of ReCoN units.
@@ 7.8. Analysis of time-multiplexed ReCoN
Accelerator design variants. In Figure 15, we show the layout of different variants of MicroScopiQ with varying number of ReCoN units, each of which can be employed based on application requirements. For area critical applications, design A, with a single ReCoN can be employed to have minimal area utilization with a slight increase in latency due to access conflicts to ReCoN. With the addition of more ReCoN units, fewer rows of PEs share a single ReCoN, dramatically reducing access conflicts to the shared ReCoN unit (design B). When the number of ReCoN units equals the number of rows, each PE row gets a dedicated ReCoN unit as shown in design C, which can be employed for latency critical applications. As we shall demonstrate subsequently, we identified that in up to a $128\times 128$ PE array size, a total of 8 ReCoN units are sufficient to achieve peak performance with zero access conflicts to ReCoN.
Optimal # of ReCoN units. For a $64\times 64$ MicroScopiQ accelerator, we study the number of access conflicts for different number of ReCoN units in Figure 16(b). Evidently, the number of access conflicts to ReCoN which is measured as the percent of total number of accesses to ReCoN that result in conflicts is under 3%. With the progressive addition of more ReCoN units, the access conflicts tend to 0%. As observed in Figure 12, despite the access conflicts, 1 ReCoN unit is sufficient to reach optimal inference performance in an iso-accuracy scenario. For latency critical applications, in Figure 18(a), we quantify the performance improvement and impact on compute area by incorporating multiple ReCoN units for a LLaMA3-8B FM. Up to 8 ReCoN units enables peak performance with 21% improved latency and only 1.58×1.58$\times$1.58 × higher compute area. For large array sizes, up to 8 ReCoN units are sufficient for peak performance.
Figure 16. (a) Annotated MicroScopiQ floorplan ($64\times 64$ array). (b) Percentage access conflicts to ReCoN for $64\times 64$ PE array.
@@ 7.9. Scalability and Timing Analysis
Scalability. To study how MicroScopiQ scales to different PE array sizes, we compare the total area (including on-chip weight and iAct/oAct buffers) against OliVe at three different scales: $8\times 8,16\times 16,128\times 128$ in Figure 17. Following (Sumbul et al., 2022), for $8\times 8$ array we employ 16 kB iAct and oAct buffers and 32 kB weight buffer. We also depict the accelerator floorplan for a $64\times 64$ array in Figure 16(a). We progressively scale the buffers for larger arrays to ensure full array utilization. Since varying the number of ReCoN units trades off area for performance, we also compare three versions of MicroScopiQ with 1, 2, 8 ReCoN units. For the single ReCoN variant of MicroScopiQ, as the design scales to higher array sizes, the area overhead of ReCoN decreases ($128\times 128$ has overhead of 3%), with the on-chip area being dominated by the large PE array and buffers. We observe similar trends for 2 or 8 ReCoN units. At large PE array sizes such as $128\times 128$, adding 8 ReCoN units results in a minimal overhead of just 11%. Particularly at large array sizes the column-wise bus from PEs to ReCoN can be routed over logic as the PE array scales. With same buffer configurations, OliVe has a higher area compared to a single ReCoN MicroScopiQ variant, while comparable area to MicroScopiQ with 8 ReCoN units. At all variants, MicroScopiQ has 23×2-3$\times$2 - 3 × higher performance than OliVe (see Table 5).
Timing Analysis. MicroScopiQ at all scales attains a peak clock frequency of 1 GHz. The critical path for MicroScopiQ is from the local weights registers to multipliers within the PE. The column-wise bus (the longest wire in MicroScopiQ) from PEs to ReCoN is not on the critical path and therefore does not limit scalability.
@@
Implementation overhead on NoC-based accelerators. The global column-wise bus is a design commonly seen in AI ASICs like (Chen et al., 2019; Firoozshahian et al., 2023; Shao et al., 2019; Tong et al., 2024). These global buses are used in conjunction with the NoC in these designs for data distribution, layout reordering (Tong et al., 2024). We modify two large-scale accelerators MTIA and Eyeriss-v2 that employ NoCs in their design to support MicroScopiQ, by incorporating ReCoN functionality in the accelerator NoCs and modifying PEs to support all MicroScopiQ specific operations. As shown in Figure 18(b), compared to the baseline design of MTIA and Eyeriss-v2, integrating ReCoN and MicroScopiQ PEs results in only a $3\%$ and $2.3\%$ increase in compute area, respectively. Since these designs already employ NoCs, integration overhead of MicroScopiQ is minimal. This demonstrates that the benefits achieved by MicroScopiQ are attained at minimal to no additional cost.